OpenStudy (anonymous):

what is the area between y=x and y=(x^2/2)

OpenStudy (anonymous):

so 4/3?

OpenStudy (raden):

yes..

OpenStudy (anonymous):

thats what i got using the integral process and it says its wrong

OpenStudy (anonymous):

can you help me with similar problems

OpenStudy (raden):

well, if u want using the integral process ... first, find the interval x's between that curves y=x^2/2 y=x x^2/2 = x x^2-2x=0 x(x-2)=0 for zeroes, satisfied for x=0 or x=2 now, calculate int(x-(x^2/2)dx from 0 to 2 = [1/2*x^2 - 1/6*x^3] from 0 to 2 = [1/2*(2)^2 - 1/6*(2)^3]-[1/2*(0)^2 - 1/6*(0)^3] = (2-8/6) - 0 = 8/6 = 4/3

OpenStudy (anonymous):

i know thats what i got using the integral process but my website for homework is marking it wrong

OpenStudy (anonymous):

what about the area between x^2+6 and 12-x^2

OpenStudy (raden):

wait still thinking was wrong with me for number one

OpenStudy (raden):

yea, i got it... 2-8/6 = 12/6 - 8/6 = 4/6 = 2/3 it should be 2/3 not 4/3

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

where did you get the 8/6 from? I got 2-4/6

OpenStudy (raden):

1/6*(2)^3 = 1/6 * 8 = 8/6, right ?

OpenStudy (anonymous):

oooooo i copied my notes down wrong

OpenStudy (anonymous):

thx a lot! can you help with similar problems...i have a bunch im stuck on

OpenStudy (raden):

for #2 do same idea like number one. first, can u find the intersection of y=x^2+6 and y=12-x^2 (*just for x) and then do integral process again :)